OpenStudy (anonymous):

solve for c 7.85 - 2.15c = 20.75

OpenStudy (anonymous):

7.85 - 2.15c = 20.75 7.85 - 7.85 - 2.15c = 20.75 - 7.85 2.15 = 12.4 c = 6 ?

OpenStudy (anonymous):

two things wrong here.

OpenStudy (anonymous):

First, you subtracted 7.85 from both sides - This is good! but, what is 20.75 - 7.85?

OpenStudy (anonymous):

12.9

OpenStudy (anonymous):

typed it wrong :s

OpenStudy (anonymous):

good; your other problem is after that, you literally skipped a step.

OpenStudy (anonymous):

so you are left with -2.15c = 12.9 after you subtract 7.85 from both sides.

OpenStudy (anonymous):

what do you do now?

OpenStudy (anonymous):

forgot the negative again. divide by - 2.15 on both sides

OpenStudy (anonymous):

mmhmm, and what does this give you for an answer?

OpenStudy (anonymous):

instead of 6, -6

OpenStudy (anonymous):

yep yep, good job ;)